(3 points) Given the fourth order homogeneous constant coefficient equation y + 10y + 9y-0 1) the auxiliary equation is ar
Answer #1

Given differential equation is:

y''''+10y''+9y=0

1) The auxiliary equation for a fourth order homogeneous constant coefficient differential equation

ay''''+by'''+cy''+dy'+ey=0 is:

ar^4+br^3+cr^2+dr+e=0

Therefore the auxiliary equation is:

r^4+10r^2+9=0

2) the roots of this equation are:

(r^2+1)(r^2+3)=0

\therefore r=\pm i, \pm 3i

3) A fundamental set of solutions is:

y_{1}=a\sin x, y_{2}=b\cos x, y_{3}=c\sin 3x, y_{4}=d\cos 3x

Therefore the general solution is:

y=a\sin x+b\cos x+c\sin 3x+d\cos 3x

4) the first three derivatives of this are:

y'=a\cos x-b\sin x+3c\cos 3x-3d\sin 3x

y''=-a\sin x-b\cos x-9c\sin 3x-9d\cos 3x

y'''=-a\cos x+b\sin x-27c\cos 3x+27d\sin 3x

Substitute given initial conditions into these.

y(0)=0, y'(0)=3, y'''(0)=-16, y'''(0)=-51

y(0)=b+d=0

y''(0)=a+3c=3

y''(0)=-b-9d=-16

y'''(0)=-a-27c=-51

By solving these, we get

a=-3, b=-2, c=2, d=2

So, the solution to the initial value problem is:

y(x)=-3\sin x-2\cos x+2\sin 3x+2\cos 3x